OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x − 2y if x = −1 and y = −4 −12 2 4 −4

OpenStudy (anonymous):

Plug and chug. 4.

OpenStudy (anonymous):

chug?

OpenStudy (anonymous):

Its just a saying :) You plug in for x as -1 and plug in for y as -4

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

i always get something else!

OpenStudy (anonymous):

4(-1)-2(-4) = -4-(-8) = -4+8 = 4